OpenStudy (anonymous):

Calculas 1help A spherical balloon is to be deflated so that its radius decreases at a constant rate of . At what rate must air be removed when the radius is

OpenStudy (paxpolaris):

when the radius is ?

OpenStudy (paxpolaris):

you want to find the rate of change of the volume:

OpenStudy (paxpolaris):

dV/dt

OpenStudy (paxpolaris):

use the chain rule: dV/dt = dV/dr * dr/dt

OpenStudy (anonymous):

constant rate of 15cm/min and radius 9cm can you show me the steps im confused

OpenStudy (paxpolaris):

dr/dt = 15 cm per minute ... given

OpenStudy (anonymous):

ok check

OpenStudy (paxpolaris):

and what't the formula for Volume of sphere?

OpenStudy (anonymous):

V=pi/3r^2h

OpenStudy (anonymous):

so the r is plugged in?

OpenStudy (paxpolaris):

SPHERE... \[V(r)=\frac 43 \pi r^3\] correct?

OpenStudy (anonymous):

oh yeahhh

OpenStudy (paxpolaris):

so dV/dr = ? (you need to get the formula for dV/dr before plugging r)

OpenStudy (anonymous):

so I derive the equation with respect to r instead of t?

OpenStudy (paxpolaris):

yes

OpenStudy (anonymous):

is dv/dr=3r^2??

OpenStudy (paxpolaris):

that times 4pi/3 ...

OpenStudy (paxpolaris):

\[dV/dr=4\pi r^2\]

OpenStudy (anonymous):

wait isn't the derivative of 4pi/3=0???

OpenStudy (paxpolaris):

the derivative of x^3 is: 3x^2 the derivative of 2x^3 is 6x^2 the derivative of ax^3 = 3ax^2

OpenStudy (anonymous):

yes I know that but how come I didn't derive 4pi/3??

OpenStudy (paxpolaris):

*... for any constant a. 4pi/3 is just a constant.

OpenStudy (anonymous):

so constants I leave alone?? and I plugged into the equation and I have dv/dr=4pi(81) is that right so far?

OpenStudy (paxpolaris):

yes just like when taking the derivative of 2x^3 ... the 2 doesn't just disappear

OpenStudy (anonymous):

how does the pi disapper I get 4860pi cm^3/min but my book answer has no pi

OpenStudy (paxpolaris):

right, dV/dr= 324 pi cm^3 per cm dr/dt = 15 cm per min dV/dt= dV/dr * dr/dt = 324pi*15 = 4860pi cm^3/min

OpenStudy (paxpolaris):

pi is just a number 3.14159...

OpenStudy (paxpolaris):

\[\Large 4860\pi \approx 15268.1403\]

OpenStudy (anonymous):

oh yeah it does have pi thank you!